# Language Inclusion for Boundedly-Ambiguous Vector Addition Systems is Decidable

### v3

@@ 1. Introduction
For each $k\in\mathbb{N}$ and a $k$-ambiguous VASS one can build in elementary time a VASS with a downward-closed set of accepting configurations which recognises the complement of its language.
@@
Organisation of the paper. In Section 2 we introduce the necessary notions. Then in Section 3 we present results concerning deterministic VASS. First, we show Theorem 1. Next, we prove that the inclusion problem of a VASS language in a language of a deterministic VASS, a $k$-deterministic VASS or a VASS with holes (to be defined) is in Ackermann. This is achieved by reducing the VASS reachability problem. In Section 4 we define adding a regular lookahead to VASS. Then we show that with a carefully chosen lookahead we can reduce the inclusion problem of a VASS language in an unambiguous VASS language into the inclusion problem of a VASS language in the language of deterministic VASS with holes. The latter is in Ackermann due to Section 3 so the former is also in Ackermann. In Section 5 we present the proof of Theorem 2 which is our most technically involved contribution. We also use the idea of a regular lookahead and the result proved in Section 3 about $k$-deterministic VASS. Finally in Section 6 we discuss the implications of our results and sketch possible future research directions.
@@ 2. Preliminaries
Basic notions. For $a,b\in\mathbb{N}$ we write $[a,b]$ to denote the set $\{a,a+1,\ldots,b-1,b\}$. For a vector $v\in\mathbb{N}^{d}$ and $i\in[1,d]$ we write $v[i]$ to denote the $i$-th coordinate of vector $v$. By $0^{d}$ we denote the vector $v\in\mathbb{N}^{d}$ with all the coordinates equal to zero. For a word $w=a_{1}\cdot\ldots\cdot a_{n}$ and $1\leq i\leq j\leq n$ we write $w[i..j]=a_{i}\cdot\ldots\cdot a_{j}$ for the infix of $w$ starting at position $i$ and ending at position $j$. We also write $w[i]=a_{i}$. For any $1\leq i\leq d$ by $e_{i}\in\mathbb{N}^{d}$ we denote the vector with all coordinates equal to zero except the $i$-th coordinate, which is equal to one. For a finite alphabet $\Sigma$ we denote $\Sigma_{\varepsilon}=\Sigma\cup\{\varepsilon\}$ the extension of $\Sigma$ by the empty word $\varepsilon$.
@@
Here we present only the proof for the complement of the upward-closed set $U$ as the case for downward-closed sets follows the same lines. Let $U=u_{1}\mathord{\uparrow}\cup u_{2}\mathord{\uparrow}\cup\ldots\cup u_{n}\mathord{\uparrow}$. Then
$\overline{U}$ $=\mathbb{N}^{d}\setminus U=\mathbb{N}^{d}\setminus(u_{1}\mathord{\uparrow}\cup u_{2}\mathord{\uparrow}\cup\ldots\cup u_{n}\mathord{\uparrow})$
$=(\mathbb{N}^{d}\setminus u_{1}\mathord{\uparrow})\cap(\mathbb{N}^{d}\setminus u_{2}\mathord{\uparrow})\cap\ldots\cap(\mathbb{N}^{d}\setminus u_{n}\mathord{\uparrow}).$
Thus in order to show that $||\overline{U}||$ is at most exponential w.r.t. $||U||$ we need to face two challenges. The first one is to show that representation of $(\mathbb{N}^{d}\setminus u\mathord{\uparrow})$ for $u\in\mathbb{N}^{d}$ is not too big w.r.t. size of $u$ and the second one is to show that the intersection of sets $(\mathbb{N}^{d}\setminus u\mathord{\uparrow})$ does not introduce too big blowup.
Let us first focus on the first challenge. Let $|u|$ be the biggest value that appear in $u$ i.e. $|u|=\max\{u[i]:i\in[1,d]\}$. We claim that if $v\in\mathbb{N}^{d}\setminus u\mathord{\uparrow}$ and $v[i]>|u|$ for $i\in[1,d]$ then $v+e_{i}\in\mathbb{N}^{d}\setminus u\mathord{\uparrow}$. Indeed, if $v\in\mathbb{N}^{d}\setminus u\mathord{\uparrow}$ then there is $j\in[1,d]$ such that $v[j]<u[j]$. Of course $i\neq j$, so $v+e_{i}\not\succeq u$ and thus $v+e_{i}\in\mathbb{N}^{d}\setminus u\mathord{\uparrow}$. But this means that if $\hat{v}\in(\mathbb{N}\cup\{\omega\})^{d}$ such that $\hat{v}\mathord{\downarrow}\subseteq\mathbb{N}^{d}\setminus u\mathord{\uparrow}$ and $\hat{v}$ is maximal (namely its entries cannot be increased without violating $\hat{v}\mathord{\downarrow}\subseteq\mathbb{N}^{d}\setminus u\mathord{\uparrow}$) then $\hat{v}\in([0,|u|]\cup\{\omega\})^{d}$. Thus there are only exponentially many possibilities for $\hat{v}$ and the representation of $\mathbb{N}^{d}\setminus u\mathord{\uparrow}$ is at most exponentially bigger than the representation of $u$.
Let us face now the second challenge. Let $\hat{v}_{1},\hat{v}_{2}\in([1,|u|]\cup\{\omega\})^{d}$. Observe that $v\in\hat{v}_{1}\mathord{\downarrow}\cap\hat{v}_{2}\mathord{\downarrow}$ if and only if $v[i]\leq\hat{v}_{1}[i]$ and $v[i]\leq\hat{v}_{2}[i]$ for all $i\in[1,d]$. But this means that if $\hat{v}\in(\mathbb{N}\cup\{\omega\})^{d}$ and $\hat{v}\mathord{\downarrow}\subseteq\hat{v}_{1}\mathord{\downarrow}\cap\hat{v}_{2}\mathord{\downarrow}$ is maximal then $\hat{v}\in([0,|u|]\cup\{\omega\})^{d}$. Thus, the representation of $\mathbb{N}^{d}\setminus U$ is also only at most exponentially bigger than the representation of $U$.
In order to compute the representation of $U$, one can simply check for all $\hat{v}\in([0,|u|]\cup\{\omega\})^{d}$ whether $\hat{v}\mathord{\downarrow}\subseteq\mathbb{N}^{d}\setminus U$.
@@
Vector Addition Systems with States. A $d$-dimensional Vector Addition System with States ($d$-VASS or simply VASS) $V$ consists of a finite alphabet $\Sigma$, a finite set of states $Q$, a finite set of transitions $T\subseteq Q\times\Sigma\times\mathbb{Z}^{d}\times Q$, a distinguished initial configuration $c_{I}\in Q\times\mathbb{N}^{d}$, and a set of distinguished final configurations $F\subseteq Q\times\mathbb{N}^{d}$. We write $V=(\Sigma,Q,T,c_{I},F)$. Sometimes we ignore some of the components in a VASS if they are not relevant, for example we write $V=(Q,T)$ if $\Sigma$, $c_{I}$, and $F$ do not matter. Configuration of a $d$-VASS is a pair $(q,v)\in Q\times\mathbb{N}^{d}$, we often write it $q(v)$ instead of $(q,v)$. We write $\textup{state}(q(v))=q$. The set of all configurations is denoted $\textup{Conf}=Q\times\mathbb{N}^{d}$. For a state $q\in Q$ and a set $U\subseteq\mathbb{N}^{d}$ we write $q(U)=\{q(u)\mid u\in U\}$. A transition $t=(p,a,u,q)\in T$ can be fired in the configuration $r(v)$ if $p=r$ and $u+v\in\mathbb{N}^{d}$. Then we write $p(v)\stackrel{{\scriptstyle t}}{{\longrightarrow}}q(u+v)$. We say that the transition $t\in T$ is over the letter $a\in\Sigma$ or the letter $a$ labels the transition $t$. We write $p(v)\stackrel{{\scriptstyle a}}{{\longrightarrow}}q(u+v)$ slightly overloading the notation, when we want to emphasise that the transition is over the letter $a$. The effect of a transition $t=(p,a,u,q)$ is the vector $u$, we write $\textup{eff}(t)=u$. The size of VASS $V$ is the total number of bits needed to represent the tuple $(\Sigma,Q,T,c_{I},F)$, we do not specify here how we represent $F$ as it may depend a lot on the form of $F$. A sequence $\rho=(c_{1},t_{1},c^{\prime}_{1}),(c_{2},t_{2},c^{\prime}_{2}),\ldots,(c_{n},t_{n},c^{\prime}_{n})\in(\textup{Conf}\times T\times\textup{Conf})^{+}$ is a run of VASS $V=(Q,T)$ if for all $i\in[1,n]$ we have $c_{i}\stackrel{{\scriptstyle t_{i}}}{{\longrightarrow}}c^{\prime}_{i}$ and for all $i\in[1,n-1]$ we have $c^{\prime}_{i}=c_{i+1}$. We write $\textup{trans}(\rho)=t_{1}\cdot\ldots\cdot t_{n}$. We extend the notion of labelling to runs; Labelling of a run $\rho$ is the concatenation of labels of its transitions. This run $\rho$ is from the configuration $c_{1}$ to the configuration $c^{\prime}_{n}$ and the configuration $c^{\prime}_{n}$ is reachable from the configuration $c_{1}$ by the run $\rho$. We write then $c_{1}\stackrel{{\scriptstyle\rho}}{{\longrightarrow}}c^{\prime}_{n}$, $c_{1}\stackrel{{\scriptstyle w}}{{\longrightarrow}}c^{\prime}_{n}$ if $w$ labels $\rho$ slightly overloading the notation or simply $c_{1}\longrightarrow c^{\prime}_{n}$ if the run $\rho$ is not relevant, we say that the run $\rho$ is over the word $w$.
@@
An example language which can be recognised by $k$-ambiguous VASS, but probably cannot be recognised by a $(k-1)$-ambiguous VASS is the following: language of words over $\{a,b\}$ with exactly $k$ letters $b$, which divides the word into $k+1$ blocks consisting of letters $a$, such that the first block is not the shortest one. A $k$-ambiguous $1$-VASS increments the counter on each letter $a$ in the first block, later guesses a block which is shortest or equal to the first block, and in that guessed block it decrements the counter. As there are at most $k$ other blocks, which may be guessed, there are at most $k$ accepting runs for each word. However, it is not clear how to accept this language by a $(k-1)$-ambiguous VASS.
@@ A VASS $V=(\Sigma,Q,T,c_{I},F)$
We provide a polynomial reduction of the problem to the emptiness problem in singleton-VASS which is in Ackermann. Let $V=(Q,T,c_{I},q_{F}(U\times D))$ be a $d$-VASS with up-atom $U\subseteq\mathbb{N}^{d_{1}}$ and down-atom $D\subseteq\mathbb{N}^{d_{2}}$ such that $d_{1}+d_{2}=d$. Let $U=u\mathord{\uparrow}$ for some $u\in\mathbb{N}^{d_{1}}$ and let $D=v\mathord{\downarrow}$ for some $v\in(\mathbb{N}\cup\{\omega\})^{d_{2}}$. Let us assume wlog of generality that $d_{2}=d_{U}+d_{B}$ such that for $i\in[1,d_{U}]$ we have $v[i]=\omega$ and for $i\in[d_{U}+1,d_{2}]$ we have $v[i]\in\mathbb{N}$. Let a $d$-VASS $V^{\prime}$ be the VASS $V$ slightly modified in the following way. First we add a new state $q^{\prime}_{F}$ and a transition $(q_{F},0^{d},q^{\prime}_{F})$. Next, for each dimension $i\in[1,d_{1}]$ we add a loop in state $q^{\prime}_{F}$ (transition from $q^{\prime}_{F}$ to $q^{\prime}_{F}$) with the effect $-e_{i}$, namely the one decreasing the dimension $i$, these are the dimensions corresponding to the up-atom $U$. Similarly for each dimension $i\in[d_{1}+1,d_{1}+d_{U}]$ we add in $q^{\prime}_{F}$ a loop with the effect $-e_{i}$, these are the unbounded dimensions corresponding to the down-atom $D$. Finally, for each dimension $i\in[d_{1}+d_{U}+1,d]$ we add in $q^{\prime}_{F}$ a loop with the effect $e_{i}$ (notice that this time we increase the counter values); these are the bounded dimensions corresponding to down-atom $D$. Let the initial configuration of $V^{\prime}$ be $c_{I}$ (the same as in V) and the set of final configurations $F^{\prime}$ of $V^{\prime}$ be the singleton set containing $q^{\prime}_{F}(u,(0^{d_{U}},v[d_{U}+1],\ldots,v[d_{U}+d_{B}]))$. Clearly $V^{\prime}$ is a singleton-VASS, so the emptiness problem for $V^{\prime}$ is in Ackermann. It is easy to see that the emptiness problems in $V$ and in $V^{\prime}$ are equivalent, which finishes the proof.
@@ 3.1. Lower bound
so $v^{\prime}$ is identical as $v$ on the first $d$ dimensions and on the last $(d+1)$-th dimension it keeps the sum of all the others. Notice that the transitions in $T^{\prime}$ are used both in $V_{1}$ and in $V_{2}$.
@@ 3.2. Upper bounds
For a VASS $V_{1}$ and a downward-VASS $V_{2}$ one can construct in polynomial time an updown-VASS $V$ such that $L(V)=L(V_{1})\cap L(V_{2})$.
@@
For a deterministic VASS, one can build in exponential time a downward VASS that recognises the complement of its language.
@@
We are ready to describe VASS $V^{\prime}=(\Sigma,Q^{\prime},T^{\prime},c^{\prime}_{I},F^{\prime})$. Basically speaking, it consists of $|T|+|\Sigma|+1$ copies of $V$. Concretely, the set of states $Q^{\prime}$ is the set of pairs $Q\times(T\cup\Sigma\cup\{-\})$. Let $c_{I}=q_{I}(v_{I})$. Then let $q^{\prime}_{I}\in Q^{\prime}$ be defined as $q^{\prime}_{I}=(q_{I},-)$ and we define the initial configuration of $V$ as $c^{\prime}_{I}=q^{\prime}_{I}(v_{I})$. The set of accepting configurations $F^{\prime}=F_{1}\cup F_{2}\cup F_{3}$ is a union of three sets $F_{i}$, each set $F_{i}$ for $i\in\{1,2,3\}$ is responsible for accepting words rejected by VASS $V$ due to the scenario (i) described above. We successively describe which transitions are added to $T^{\prime}$ and which configurations are added to $F^{\prime}$ in order to appropriately handle various scenarios.
@@
Now we describe the second part of $V^{\prime}$ which is responsible for words rejected by $V$ because of the scenario (2). The idea is to guess when the run over $w$ is finished. For each transition $t=(p,a,v,q)\in T$ we add $(p^{\prime},a,0^{d},q^{\prime})$ to $T^{\prime}$ where $p^{\prime}=(p,-)$ and $q^{\prime}=(q,t)$. The idea is that the run reaches the configuration in which the transition $t$ cannot be fired. Now we have to check that our guess is correct. In the state $(q,t)$ for $t\in T$ no transition changes the configuration. Namely for each $q^{\prime}=(q,t)\in Q\times T$ and each $a\in\Sigma$ we add to $T^{\prime}$ transition $(q^{\prime},a,0^{d},q^{\prime})$. We now add to $F^{\prime}$ the set $F_{2}=\{(q,t)(v)\mid v+\textup{eff}(t)\not\in\mathbb{N}^{d}\}$. Notice that $F_{2}$ can be easily represented as a polynomial union of down-atoms. It is easy to see that indeed $V^{\prime}$ accepts by $F_{2}$ exactly words $w$ such that there is a run of $V$ over some prefix $v$ of $w$ but reading the next letter would decrease one of the counters below zero.
The last part of $V^{\prime}$ is responsible for the words $w$ rejected by $V$ because of the scenario (3), that is, $w$ has a prefix $va$ such that there is a run over $v\in\Sigma^{*}$ in $V$ but then in the state of the reached configuration there is no transition over the letter $a\in\Sigma$. To accept such words for each state $p\in Q$ and letter $a\in\Sigma$ such that there is no transition of a form $(p,a,v,q)\in T$ for any $v\in\mathbb{N}^{d}$ and $q\in Q$ we add to $T^{\prime}$ transition $((p,-),a,0^{d},(p,a))$. In each state $p^{\prime}=(p,a)\in Q\times\Sigma$, we have a transition $(p^{\prime},b,0^{d},p^{\prime})$ for each $b\in\Sigma$. We also add to $F^{\prime}$ the set $F_{3}=\{(p,a)(v)\mid v\in\mathbb{N}^{d}\text{ and there is no }(p,a,u,q)\in T\text{ for }u\in\mathbb{N}^{d}\text{ and }q\in Q\}$. The size of $F_{3}$ is polynomial w.r.t. $T$.
@@
For each HVASS one can compute in exponential time a language equivalent $\varepsilon$-VASS.
@@
For a deterministic HVASS one can compute in exponential time a downward-$\varepsilon$-VASS which recognises the complement of its language.
@@
Boundedly-deterministic VASS. We define here a generalisation of a deterministic VASS, namely a $k$-deterministic VASS for $k\in\mathbb{N}$. Such VASS are later used as a tool for deriving results about $k$-ambiguous VASS in Section 5.
We say that a finite automaton $\mathcal{A}=(\Sigma,Q,T,q_{I},F)$ is $k$-deterministic if for each word $w\in\Sigma^{*}$ there are at most $k$ maximal runs over $w$. We call a run $\rho$ a maximal run over $w$ if either (1) it is a run over $w$ or (2) $w=uav$ for $u,v\in\Sigma^{*}$, $a\in\Sigma$ such that the run $\rho$ is over the prefix $u$ of $w$ but there is no possible way of extending $\rho$ by any transition labelled with the letter $a\in\Sigma$. Let us emphasise that here we count runs in a subtle way. We do not count only the maximal number of active runs throughout the word, but the total number of different runs that have ever been started during the word. To illustrate the difference better let us consider an example finite automaton $\mathcal{A}$ over $\Sigma=\{a,b\}$ with two states $p,q$ and with three transitions: $(p,a,p)$, $(p,a,q)$ and $(q,b,q)$. Then $\mathcal{A}$ has $n+1$ maximal runs over the word $a^{n}$ although only two of these runs actually survive till the end of the input word. So $\mathcal{A}$ is not $2$-deterministic even though for each input word it has at most two runs. We say that a VASS $V=(\Sigma,Q,T,c_{I},F)$ is $k$-deterministic if its control automaton is $k$-deterministic.

### v4

@@ 1. Introduction
There is an algorithm that for each $k\in\mathbb{N}$ and a $k$-ambiguous VASS $V$ constructs, in elementary time, a VASS with a downward-closed set of accepting configurations which recognises the complement of the language of $V$.
@@
Organisation of the paper. In Section 2 we introduce the necessary notions. Then in Section 3 we present results concerning deterministic VASS. First, we show Theorem 1. Next, we prove that the inclusion problem of a VASS language in a language of a deterministic VASS, a $k$-control-deterministic VASS or a VASS with holes (to be defined) is in Ackermann. This is achieved by reducing to the VASS reachability problem. In Section 4 we define adding a regular lookahead to VASS. Then we show that with a carefully chosen lookahead we can reduce the inclusion problem of a VASS language in an unambiguous VASS language into the inclusion problem of a VASS language in the language of deterministic VASS with holes. The latter is in Ackermann due to Section 3 so the former is also in Ackermann. In Section 5 we present the proof of Theorem 2 which is our most technically involved contribution. We also use the idea of a regular lookahead and the result proved in Section 3 about $k$-control-deterministic VASS. Finally in Section 6 we discuss the implications of our results and sketch possible future research directions.
@@ 2. Preliminaries
Basic notions. By $\mathbb{N}$ and $\mathbb{Z}$ we denote natural and integer numbers, respectively. For $a,b\in\mathbb{N}$ we write $[a,b]$ to denote the set $\{a,a+1,\ldots,b-1,b\}$. For a vector $v\in\mathbb{N}^{d}$ and $i\in[1,d]$ we write $v[i]$ to denote the $i$-th coordinate of vector $v$. By $0^{d}$ we denote the vector $v\in\mathbb{N}^{d}$ with all the coordinates equal to zero. For a word $w=a_{1}\cdot\ldots\cdot a_{n}$ and $1\leq i\leq j\leq n$ we write $w[i..j]=a_{i}\cdot\ldots\cdot a_{j}$ for the infix of $w$ starting at position $i$ and ending at position $j$. We also write $w[i]=a_{i}$. For any $1\leq i\leq d$ by $e_{i}\in\mathbb{N}^{d}$ we denote the vector with all coordinates equal to zero except the $i$-th coordinate, which is equal to one. For a finite alphabet $\Sigma$ we denote $\Sigma_{\varepsilon}=\Sigma\cup\{\varepsilon\}$ the extension of $\Sigma$ by the empty word $\varepsilon$.
@@
Here we present only the proof for the complement of the upward-closed set $U$ as the case for downward-closed sets follows the same lines. Let $U=u_{1}\mathord{\uparrow}\cup u_{2}\mathord{\uparrow}\cup\ldots\cup u_{n}\mathord{\uparrow}$. The complement of $U$ is a downward closed set, that can be represented as a union of down-atoms. As each down-atom is represented as a downward closure of a single vector over $\mathbb{N}\cup\{\omega\}$ it is sufficient to calculate these vectors. Each vector $x$ representing a down-atom has two properties:
  1. (1)
    it is in the “complement“ i.e. for every $i\in\{1,\ldots,n\}$ there is a coordinate $f(i)\in\{1,\ldots,d\}$ such that $x[f(i)]<u_{i}[f(i)]$.
  2. (2)
    it is maximal, i.e. for every $y$ that satisfies the first property holds $x\not\preceq y$.
Intuitively, function $f$ witnesses why $x$ is not in $U$.
Thus, for every function $f:\{1,\ldots,n\}\xrightarrow{}\{1,\ldots,d\}$ we can define one down-atom represented by some vector $x_{f}$ as the maximal vector, for which function $f$ witnesses that $x_{f}\not\in U$. For a coordinate $i\in[1,d]$ we need that $x_{f}[i]<u_{j}[i]$ for each $j\in[1,n]$ such that $j=f(i)$. Therefore
$x_{f}[i]=\min\{u_{j}[i]-1\mid j\in[1,n],f(j)=i\},$
where we put $\min(\emptyset)=\omega.$ Notice that the union of $x_{f}\downarrow$ is exactly the complement of $U$.
Observe that there is at most exponentially many different functions $f\in\{1\ldots n\}\xrightarrow{}\{1,\ldots,d\}$ and numbers that appear in descriptions of vectors $x_{f}$ are $\omega$ or are smaller than the maximal number that appear in the description of $U$. Therefore, we conclude that the proposition is true.
@@
Vector Addition Systems with States. A $d$-dimensional Vector Addition System with States ($d$-VASS or simply VASS) $V$ consists of a finite alphabet $\Sigma$, a finite set of states $Q$, a finite set of transitions $T\subseteq Q\times\Sigma\times\mathbb{Z}^{d}\times Q$, a distinguished initial configuration $c_{I}\in Q\times\mathbb{N}^{d}$, and a set of distinguished final configurations $F\subseteq Q\times\mathbb{N}^{d}$. We write $V=(\Sigma,Q,T,c_{I},F)$. Sometimes we ignore some of the components in a VASS if they are not relevant, for example we write $V=(Q,T)$ if $\Sigma$, $c_{I}$, and $F$ do not matter. A configuration of a $d$-VASS is a pair $(q,v)\in Q\times\mathbb{N}^{d}$, we often write it $q(v)$ instead of $(q,v)$. We write $\textup{state}(q(v))=q$. The set of all configurations is denoted $\textup{Conf}=Q\times\mathbb{N}^{d}$. For a state $q\in Q$ and a set $U\subseteq\mathbb{N}^{d}$ we write $q(U)=\{q(u)\mid u\in U\}$. A transition $t=(p,a,u,q)\in T$ can be fired in the configuration $r(v)$ if $p=r$ and $u+v\in\mathbb{N}^{d}$. Then we write $p(v)\stackrel{{\scriptstyle t}}{{\longrightarrow}}q(u+v)$. We say that the transition $t\in T$ is over the letter $a\in\Sigma$ or the letter $a$ labels the transition $t$. We write $p(v)\stackrel{{\scriptstyle a}}{{\longrightarrow}}q(u+v)$ slightly overloading the notation, when we want to emphasise that the transition is over the letter $a$. The effect of a transition $t=(p,a,u,q)$ is the vector $u$, we write $\textup{eff}(t)=u$. The size of VASS $V$ is the total number of bits needed to represent the tuple $(\Sigma,Q,T,c_{I},F)$, we do not specify here how we represent $F$ as it may depend a lot on the form of $F$. A sequence $\rho=(c_{1},t_{1},c^{\prime}_{1}),(c_{2},t_{2},c^{\prime}_{2}),\ldots,(c_{n},t_{n},c^{\prime}_{n})\in(\textup{Conf}\times T\times\textup{Conf})^{*}$ is a run of VASS $V=(Q,T)$ if for all $i\in[1,n]$ we have $c_{i}\stackrel{{\scriptstyle t_{i}}}{{\longrightarrow}}c^{\prime}_{i}$ and for all $i\in[1,n-1]$ we have $c^{\prime}_{i}=c_{i+1}$. We write $\textup{trans}(\rho)=t_{1}\cdot\ldots\cdot t_{n}$. We extend the notion of labelling to runs; Labelling of a run $\rho$ is a word $w$ which is the concatenation of labels of transitions of $\rho$, we also say that the run $\rho$ is over the word $w$. This run $\rho$, if not empty, is from the configuration $c_{1}$ to the configuration $c^{\prime}_{n}$ and the configuration $c^{\prime}_{n}$ is reachable from the configuration $c_{1}$ by the run $\rho$. We write then $c_{1}\stackrel{{\scriptstyle\rho}}{{\longrightarrow}}c^{\prime}_{n}$, $c_{1}\stackrel{{\scriptstyle w}}{{\longrightarrow}}c^{\prime}_{n}$ if $w$ labels $\rho$ slightly overloading the notation or simply $c_{1}\longrightarrow c^{\prime}_{n}$ if the run $\rho$ is not relevant. The empty run can go from any configuration to itself, and it is labeled with $\varepsilon$.
@@
An example language which can be recognised by $k$-ambiguous VASS, but probably cannot be recognised by a $(k-1)$-ambiguous VASS is the following: language of words over $\{a,b\}$ with exactly $k$ letters $b$, which divides the word into $k+1$ blocks consisting of letters $a$, such that the first block is not the shortest one.block it decrements the counter. A $k$-ambiguous $1$-VASS increments the counter on each letter $a$ in the first block, later guesses a block which is the shortest one, and in that guessed block it decrements the counter. As there are at most $k$ other blocks, which may be guessed, there are at most $k$ accepting runs for each word. However, it is not clear how to accept this language by a $(k-1)$-ambiguous VASS. The problematic scenario is if the first block is the longest one.
@@ A VASS $V=(\Sigma,Q,T,c_{I},F)$
We provide a polynomial reduction of the problem to the emptiness problem in singleton-VASS which is in Ackermann. Let $V=(Q,T,c_{I},q_{F}(U\times D))$ be a $d$-VASS with up-atom $U\subseteq\mathbb{N}^{d_{1}}$ and down-atom $D\subseteq\mathbb{N}^{d_{2}}$ such that $d_{1}+d_{2}=d$. Let $U=u\mathord{\uparrow}$ for some $u\in\mathbb{N}^{d_{1}}$ and let $D=v\mathord{\downarrow}$ for some $v\in(\mathbb{N}\cup\{\omega\})^{d_{2}}$. Let us assume wlog of generality that $d_{2}=d_{U}+d_{B}$ such that for $i\in[1,d_{U}]$ we have $v[i]=\omega$ and for $i\in[d_{U}+1,d_{2}]$ we have $v[i]\in\mathbb{N}$. Let a $d$-VASS $V^{\prime}$ be the VASS $V$ slightly modified in the following way. First we add a new state $q^{\prime}_{F}$ and a transition $(q_{F},0^{d},q^{\prime}_{F})$. Next, for each coordinate $i\in[1,d_{1}]$ we add a loop in state $q^{\prime}_{F}$ (transition from $q^{\prime}_{F}$ to $q^{\prime}_{F}$) with the effect $-e_{i}$, namely the one decreasing the coordinate $i$, these are the coordinates corresponding to the up-atom $U$. Similarly for each coordinate $i\in[d_{1}+1,d_{1}+d_{U}]$ we add in $q^{\prime}_{F}$ a loop with the effect $-e_{i}$, these are the unbounded coordinates corresponding to the down-atom $D$. Finally, for each coordinate $i\in[d_{1}+d_{U}+1,d]$ we add in $q^{\prime}_{F}$ a loop with the effect $e_{i}$ (notice that this time we increase the counter values); these are the bounded coordinates corresponding to down-atom $D$. Let the initial configuration of $V^{\prime}$ be $c_{I}$ (the same as in V) and the set of final configurations $F^{\prime}$ of $V^{\prime}$ be the singleton set containing $q^{\prime}_{F}(u,(0^{d_{U}},v[d_{U}+1],\ldots,v[d_{U}+d_{B}]))$. Clearly $V^{\prime}$ is a singleton-VASS, so the emptiness problem for $V^{\prime}$ is in Ackermann. It is easy to see that the emptiness problems in $V$ and in $V^{\prime}$ are equivalent, which finishes the proof.
@@ 3.1. Lower bound
so $v^{\prime}$ is identical as $v$ on the first $d$ coordinates and on the last $(d+1)$-th coordinate it keeps the sum of all the others. Notice that the transitions in $T^{\prime}$ are used both in $V_{1}$ and in $V_{2}$.
@@ 3.2. Upper bounds
There is an algorithm that for a VASS $V_{1}$ and a downward-VASS $V_{2}$ constructs in polynomial time an updown-VASS $V$ such that $L(V)=L(V_{1})\cap L(V_{2})$.
@@
There is algorithm, that for a deterministic VASS $V$ constructs, in exponential time, a downward VASS that recognises the complement of the language of $V$.
@@
We are ready to describe VASS $V^{\prime}=(\Sigma,Q^{\prime},T^{\prime},c^{\prime}_{I},F^{\prime})$. Basically speaking, it consists of $12\cdot|\Sigma|+$ 1 copies of $V$. Concretely, the set of states $Q^{\prime}$ is the set of pairs $Q\times(\Sigma\times\{2,3\}\cup\{-\})$. Let $c_{I}=q_{I}(v_{I})$. Then let $q^{\prime}_{I}\in Q^{\prime}$ be defined as $q^{\prime}_{I}=(q_{I},-)$ and we define the initial configuration of $V$ as $c^{\prime}_{I}=q^{\prime}_{I}(v_{I})$. The set of accepting configurations $F^{\prime}=F_{1}\cup F_{2}\cup F_{3}$ is a union of three sets $F_{i}$, each set $F_{i}$ for $i\in\{1,2,3\}$ is responsible for accepting words rejected by VASS $V$ due to the scenario (i) described above. We successively describe which transitions are added to $T^{\prime}$ and which configurations are added to $F^{\prime}$ in order to appropriately handle various scenarios.
@@
Now we describe the second part of $V^{\prime}$ which is responsible for words rejected by $V$ because of the scenario (2). The idea is to guess when the run over $w$ is finished. For each label $a\in\Sigma$ we add $(p^{\prime},a,0^{d},p^{\prime\prime})$ to $T^{\prime}$ where $p^{\prime}=(p,-)$ and $q^{\prime\prime}=(p,(a,2))$. The idea is that the run reaches the configuration in which the transition labeled with $a$ cannot be fired. Now we have to check that our guess is correct. In the state $(p,(a,2))$ for $t\in T$ no transition changes the configuration. Namely for each $p^{\prime\prime}=(p,(a,2))\in Q\times\Sigma$ and each $b\in\Sigma$ we add to $T^{\prime}$ a transition $(p^{\prime\prime},b,0^{d},p^{\prime\prime})$. We now add to $F^{\prime}$ the set $F_{2}=\{(p,(a,2))(v)\mid v+\textup{eff}(t)\not\in\mathbb{N}^{d}\text{ for all }t=(p,a,\textup{eff}(t),q)\in T\text{ where }q\in Q\}$. Notice that $F_{2}$ can be easily represented as a polynomial union of down-atoms. It is easy to see that indeed $V^{\prime}$ accepts by $F_{2}$ exactly words $w$ such that there is a run of $V$ over some prefix $v$ of $w$ but reading the next letter would decrease one of the counters below zero.
The last part of $V^{\prime}$ is responsible for the words $w$ rejected by $V$ because of the scenario (3), that is, $w$ has a prefix $va$ such that there is a run over $v\in\Sigma^{*}$ in $V$ but then in the state of the reached configuration there is no transition over the letter $a\in\Sigma$. To accept such words for each state $p\in Q$ and letter $a\in\Sigma$ such that there is no transition of a form $(p,a,v,q)\in T$ for any $v\in\mathbb{Z}^{d}$ and $q\in Q$ we add to $T^{\prime}$ transition $((p,-),a,0^{d},(p,(a,3)))$. In each state $p^{\prime}=(p,(a,3))\in Q\times\Sigma$, we have a transition $(p^{\prime},b,0^{d},p^{\prime})$ for each $b\in\Sigma$. We also add to $F^{\prime}$ the set $F_{3}=\{(p,(a,3))(v)\mid v\in\mathbb{N}^{d}\text{ and there is no }(p,a,u,q)\in T\text{ for }u\in\mathbb{Z}^{d}\text{ and }q\in Q\}$. The size of $F_{3}$ is polynomial w.r.t. $T$.
@@
There is an algorithm, that for a given HVASS computes, in exponential time, a language equivalent $\varepsilon$-VASS.
@@
There is an algorithm, that for a deterministic HVASS $V$ computes, in exponential time, a downward-$\varepsilon$-VASS which recognises the complement of the language of $V$.
@@
Boundedly-control-deterministic VASS. We define here a generalisation of a deterministic VASS, namely a $k$-control-deterministic VASS for $k\in\mathbb{N}$. Such VASS are later used as a tool for deriving results about $k$-ambiguous VASS in Section 5.
We say that a finite automaton $\mathcal{A}=(\Sigma,Q,T,q_{I},F)$ is $k$-deterministic if for each word $w\in\Sigma^{*}$ there are at most $k$ maximal runs over $w$. We call a run $\rho$ a maximal run over $w$ if either (1) it is a run over $w$ or (2) $w=uav$ for $u,v\in\Sigma^{*}$, $a\in\Sigma$ such that the run $\rho$ is over the prefix $u$ of $w$ but there is no possible way of extending $\rho$ by any transition labelled with the letter $a\in\Sigma$. Let us emphasise that here we count runs in a subtle way. We do not count only the maximal number of active runs throughout the word, but the total number of different runs that have ever been started during the word. To illustrate the difference better let us consider an example finite automaton $\mathcal{A}$ over $\Sigma=\{a,b\}$ with two states $p,q$ and with three transitions: $(p,a,p)$, $(p,a,q)$ and $(q,b,q)$. Then $\mathcal{A}$ has $n+1$ maximal runs over the word $a^{n}$ although only two of these runs actually survive till the end of the input word. So $\mathcal{A}$ is not $2$-deterministic even though for each input word it has at most two runs. We say that a VASS $V=(\Sigma,Q,T,c_{I},F)$ is $k$-control-deterministic if its control automaton is $k$-deterministic.